# Conceptualization, Methodology, Software, Writing

### v1

@@ 3.2 Feature Collapsing
Feature collapsing is a commonly observed occurrence within ViT architectures. It refers to the phenomenon where the features extracted from different image patches lose their distinctiveness and become increasingly similar or indistinguishable as the network depth grows. This phenomenon primarily arises due to the nature of the attention mechanism used in ViT, which progressively aggregates information from various image patches as it traverses through the network layers. Such a process can be mathematically described by representing Eq. (5) as a weighted summation of feature patches in the following manner:
$\tilde{X}^{i}_{l,h}=\sum^{N}_{j=1}A^{i,j}_{l,h}X^{j}_{l,h},\text{subject to}\sum^{N}_{j=1}A^{i,j}_{l,h}=1,\text{for}\ i=[1,2,...,N].$ (7)
Here, $\tilde{X}^{i}_{l,h}$ denotes the feature vector originating from patch $i$ and represents a weighted average of features from all patches. The weights are determined by the values in the attention map $A_{l,h}$. Although the attention mechanism aims to capture global relationships between different feature patches [7], it can lead to the loss of feature diversity, resulting in feature collapsing. In order to provide evidence for this observation, Fig. 3 presents a visual representation of the feature similarity matrix for the ViT. This matrix is constructed by calculating the cosine similarity between distinct feature patches extracted by the ViT model. The key insight revealed by the figure is that, as the features progress from the shallower layers of the model to the deeper layers, there is a noticeable trend towards increased similarity among them.
@@
In this equation, the identity projection $X_{l}$ runs in parallel to the MHSA operation. Intuitively, because the features in $X_{l}$ exhibit greater diversity before passing through the MHSA operation, the summation of the two feature vectors preserves the characteristics from the previous layer. Consequently, the output retains more diverse features. Empirical evidence supporting this phenomenon can be found in [32]. Drawing inspiration from these discoveries, this study introduces a novel approach to tackling this issue. Specifically, it suggests an innovative alternative residual connection between attention layers (without bypassing the attention mechanism) designed to curb the rapid expansion of the attention across the head dimension, which is the main source of feature collapsing, and increase feature diversity as shown in Fig. 3.
@@ 3.3 Residual Attention
where $\alpha\in]0;1[$. After obtaining the balanced attention scores $S_{l}$, following Eq. (2), the softmax function is applied to compute the attention weights $A_{l}$ required for obtaining the new scaled dot-product attention output, as depicted in Fig. 2(b). Given the implementation of residual attention, it can be easily included in both existing single-scale or multi-scale vision transformer architectures, maintaining a similar computational cost.
@@ 3.4 Attention Globality
Attention globality is a concept that reflects the capacity of the attention mechanism to encompass and consider information from all patches across an image rather than being restricted to local or nearby elements. This capability becomes particularly pronounced in the deeper layers of a ViT because these layers can extract high-level representations that account for long-range relationships between patches. To quantify the globality of the attention mechanism across both heads and layers, researchers in [8] introduce a metric known as the non-locality metric. This metric calculates, for each query patch $i$, the relative positional distances to all key patches $j$ weighted by their attention scores $A^{i,j}_{l,h}$. The resulting sum is then averaged by the number of patches to derive the non-locality metric for a specific head $h$. This can further be averaged across attention heads to obtain the non-locality metric for the entire layer. Mathematically, this metric is defined as follows:
[…]
Here, $\delta^{i,j}||$ represents the relative positional distance between query patch $i$ and key patch $j$, $D_{l,h}$ denotes the non-locality metric within a given layer $l$ and head $h$, while $D_{l}$ represents $D_{l,h}$ averaged along the head dimension. Moreover, since $\delta^{i,j}||$ remains constant throughout the network (i.e., the relative positions of patches do not change), the value of $D_{l,h}$ relies entirely on $A_{l,h}$. Furthermore, as the non-locality metric $D$ increases in deeper layers (indicating a shift towards global attention), it suggests that the attention matrix $A_{l,h}$ (consequently $A_{l}$) in these layers generally has a greater capacity to capture global relationships across the different attention heads compared to attention matrix from the previous layer. This property of the attention mechanism can be harnessed to demonstrate the advantages of employing residual attention in slowing down the process of globalizing $A_{l}$. Taking into account Eq. (10), which defines the calculation of $A_{l}$ with a residual connection, it considers attention scores not only from the current layer $l$ but also from previous layers. Consequently, the globality of layer $l$ in ReViT, at a given head, can be represented as:
[…]
@@ 4 Experiments
In this section, we present the implementation details of the proposed method and tests conducted on some challenging datasets for image classification, object detection, and instance segmentation. In particular, we start by giving the implementation details of our method in various settings. Subsequently, we report the test results on five image classification benchmarks, namely ImageNet1K, CIFAR-10, CIFAR-100, Oxford Flowers-102, and Oxford-IIIT Pet. To proceed, the test results under position and scale changes (i.e., translation invariance) on Oxford Flowers-102 and Oxford-IIIT Pet datasets are reported. Such datasets were chosen for translation invariance testing due to their higher quality, which helps preserve important information even after applying translation transformations like scale reduction. Moreover, we investigate the attention mechanism of the networks trained on ImageNet1K, both quantitatively using the non-locality metric defined in Eq. (11-12) and qualitatively comparing the learned feature maps extracted with the GradCAM algorithm between the proposed approach and the standard ViT. Furthermore, we report the test results that show the effectiveness of our method when implemented on top of spatial-aware multiscale architectures like Swin and MViTv2 on downstream tasks like object detection and instance segmentation on the COCO2017 dataset.
@@ 4.1 Implementation Details
The proposed ReViT method is implemented as an augmented version of the standard ViT architecture. Specifically, we integrate the residual attention module on top of the standard ViT implementation using the PyToch framework. Regarding the network version used in this study, we only relied on the base version of ViT containing 12 layers and denominated as ViT-B. To this end, we name our network ReViT-B. Additionally, as mentioned above, we conduct experiments that test the effectiveness of our method applied in the multiscale architectures, namely MViTv2 and Swin. To implement such networks, we start from their official PyTorch implementations and extend them with our module. Moreover, for MViTv2 and Swin, we exploit their tiny versions, i.e., MViTv2-T and Swin-T with 12 layers each, and name our implementations ReMViTv2-T and ReSwin-T. The rationale behind using the smaller versions of each network is that they require less computational resources, and usually, the bigger versions require pre-training on large-scale datasets like ImageNet21K or even bigger ones.
To train our networks, we use the same configurations reported in their official studies and run the experiments on an NVIDIA V100 GPU with 32GB of RAM. Specifically, for image classification on ImageNet1K, all models were trained for $300$ epochs, with an input image size of $224\times 224$, using gradient clipping and cosine scheduling learning rate warmup/decay. For the ReViT-B network training, Adam [17] was used as optimizer with a base learning rate set to $0.001$, $30$ warmup epochs, an effective batch size of $4096$, and the weight decay set to $0.3$. Instead, for ReSwin-T model training, AdamW optimizer [26] was used by setting a base learning rate to $0.001$, 2020 warmup epochs, an effective batch size of $1024$, and the weight decay set to $0.05$. Finally, the remaining ReMViTv2-T network training was performed using the AdamW optimizer with a base learning rate set to $0.002$, $70$ warmup epochs, an effective batch size of $2048$, and the weight decay set to $0.1$. Regarding other image classification tasks, we use the same training setups reported in [7] for a fair comparison. Moreover, regarding the object detection and instance segmentation tasks, since such tasks require local awareness of features, the experiments were performed only using ReSwin-T and ReMViTv2-T. These models were trained for $36$ epochs using AdamW as an optimizer and linear scheduling learning rate decay. For ReSwin-T network training, the initial learning rate was set to $0.0001$, with a weight decay of $0.05$, and a batch size set to $16$. Instead, for ReMViTv2-T model training, the initial learning rate was set to $0.00016$, with a weight decay of $0.1$, and a batch size set to $64$.
@@ 4.2 Image classification
Subsequently, ReViT-B was tested on multiple image classification benchmarks, including CIFAR-10, CIFAR-100, Oxford Flowers-102, and Oxford-IIIT Pet. The performance of ReViT-B is then compared with that of the ViT-B model, and the results are presented in Table 3. Notably, ReViT-B, with the residual attention, consistently outperforms ViT-B across all datasets with a minimum improvement of 0.6% in CIFAR10 and a maximum of 3.4% in CIFAR100. This improvement can be attributed to the residual attention module’s ability to capture a wider range of visual features, thus enhancing the network’s capability to distinguish between different classes based on the image content. The other models were not tested on these datasets because their official studies do not provide experimental results and focus only on large-scale datasets.
@@
Model CIFAR-10 CIFAR-100 Oxford-Flowers 102 Oxford-IIIT Pet
ViT-B [7] 98.1 87.1 89.5 93.8
ReViT-B 98.7 90.5 91.1 94.9
@@ 4.3 Translation Invariance
In this study, the datasets used for visual recognition typically feature the object of interest centered within the image. This is in stark contrast to real-world scenarios where objects can be situated anywhere within the scene. To assess ReViT robustness, two experiments were devised to simulate varying scales and positions of the target object within the images as shown in Fig. 5. The performance of ReViT-B and ViT-B on the Oxford Flowers-102 and Oxford-IIIT Pet datasets are compared in these experiments.
The first experiment evaluates the models’ ability to handle horizontal and vertical translations. To achieve this, the input images are horizontally and vertically shifted on the image plane by {15, 30, 45, 60}% of pixels, with the empty areas filled with zeros to maintain the same input image size as shown in Fig. 5(a) and (b). As indicated in Table 4 and Fig. 6(a), both models exhibit a similar and impressive level of horizontal translation invariance, with only a minor reduction in performance. Overall, ReViT-B demonstrates slightly superior horizontal translation invariance, especially at the 60% translation level. For the Oxford Flowers-102 dataset, ReViT-B performance decreases by 2.9% less than ViT-B at a 60% translation, while for the Oxford-IIIT Pet dataset, the decrease is 1.0% less. Similar trends are observed for vertical translation invariance, as illustrated in Table 5 and Fig. 6(b). Although both models show a comparable drop in performance, ReViT-B displays better overall vertical translation invariance, particularly on the Oxford Flowers-102 dataset, where its performance decreases by 3.6% less than ViT-B.
In the second experiment, the scale invariance of both models is evaluated, and the results are presented in Table 6 and Fig. 6(c). To generate images at different scales, the original images are resized by reducing their width and height by {15, 30, 45, 60}%, with zero-padding to maintain the correct input size as illustrated in Fig. 5(c). Unlike the first experiment, ReViT-B consistently outperforms ViT-B in terms of scale invariance. In the case of the Oxford Flowers-102 dataset, ReViT-B shows up to 19.5% less reduction in performance accuracy, while for the Oxford-IIIT Pet dataset, it exhibits up to 4.7% less decrease than ViT-B. These results highlight that ReViT-B possesses significantly better scale invariance compared to ViT-B.
@@ 4.4 Residual Attention and Globality
The investigation into the global reach of MHSA within vision transformers, both with and without residual attention, aims to examine the hypothesis that residual attention decelerates the globalization of the attention mechanism. To assess this hypothesis quantitatively, we employ the non-locality metric calculated according to Eq. (11). Additionally, the metric is computed at both the head and layer levels. The results of this analysis, depicted in Fig. 7, display the average metric across 256 input images from the ImageNet1K validation set for ViT-B and ReViT-B models trained on the same dataset. The x-axis on the plots corresponds to the layer index, while the mean non-locality for each head or layer is represented on the y-axis. Fig. 7(a) illustrates that in ViT-B, the non-locality of each head increases as it delves deeper into the layers, approaching globality which corresponds to the findings in [7, 8]. A similar trend is observed in ReViT-B, as seen in Fig. 7(b), albeit with some noteworthy distinctions. Notably, ViT-B exhibits a more consistent and monotonic increase in non-locality, with minimal variations between different heads, particularly after layer 5. In contrast, ReViT-B displays more diverse differences between heads, with some heads finding difficulty in escaping the locality while others become completely global, resulting in a less monotonic increase in non-locality. In summary, the results substantiate the hypothesis that the attention receptive field of heads in vision transformers becomes global in deeper layers, a phenomenon observed in both ViT-B and ReViT-B models. However, these findings also reveal that ReViT-B retains the ability to extract global relationships in deeper layers while still maintaining some heads that remain localized with relatively low non-locality values. Simultaneously, ViT-B is confined solely to global relationships, leading to feature collapse. To provide a broader perspective, Fig. 7(c) presents the non-locality of ViT-B and ReViT-B at the layer level, represented as the average non-locality metric across all heads within each layer. It is evident that when computed using the same 256 input examples, ReViT-B consistently maintains lower non-locality values (except the first), indicating its ability to incorporate low-level features, with the help of relatively local heads, into the representations learned from heads that exhibit non-locality escape (i.e., become global).

### v2

@@ 3.2 Feature Collapsing
Feature collapsing is a commonly observed occurrence within ViT architectures. It refers to the phenomenon where the features extracted from different image patches lose their distinctiveness and become increasingly similar or indistinguishable as the network depth grows. This phenomenon primarily arises due to the nature of the attention mechanism used in ViT, which progressively aggregates information between image patches as it traverses through the network layers. Such a process can be mathematically described by representing Eq. (4) as a weighted summation of feature patches in the following manner:
$\tilde{X}^{i}_{l,h}=\sum^{N}_{j=1}A^{i,j}_{l,h}V^{j}_{l,h},\text{subject to}\sum^{N}_{j=1}A^{i,j}_{l,h}=1,\text{for}\ i=[1,2,...,N].$ (7)
Here, $\tilde{X}^{i}_{l,h}$ denotes the feature vector originating from patch $i$ and represents a weighted average of features from all patches j. The weights are determined by the values in the attention map $A_{l,h}$. Although the attention mechanism aims to capture global relationships between different feature patches [6], it can lead to the loss of feature diversity, resulting in feature collapsing. In order to provide evidence for this observation, Fig. 3 presents a visual representation of the feature similarity matrix for the ViT. This matrix is constructed by calculating the cosine similarity between distinct feature patches extracted by the ViT model. The key insight revealed by the figure is that, as the features progress from the shallower layers of the model to the deeper layers, there is a noticeable trend towards increased similarity among them.
@@
In this equation, the identity projection $X_{l}$ runs in parallel to the MHSA operation. Intuitively, because the features in $X_{l}$ exhibit greater diversity before passing through the MHSA operation, the summation of the two feature vectors preserves the characteristics from the previous layer. Consequently, the output retains more diverse features. However, the residual connection between features ${X_{l}}$ proves insufficient. Empirical evidence supports this phenomenon in [35]. Drawing inspiration from these discoveries, this study introduces a novel approach to tackling this issue. Specifically, it suggests an innovative alternative residual connection between attention layers (without bypassing the attention mechanism) designed to curb the rapid expansion of the attention across the head dimension, which is the main source of feature collapsing, and increase feature diversity as shown in Fig. 3.
@@ 3.3 Residual Attention
where $\alpha\in[0,1]$. According to this definition, for $l>0$, $S_{l}$ is obtained as a sum between the scaled dot product of ($Q_{l}$ $K_{l}$) and $S_{l-1}$ weighted by $\alpha$ and $1-\alpha$ respectively. After obtaining the balanced attention scores $S_{l}$, following Eq. (2), the softmax function is applied to compute the attention weights $A_{l}$ required for obtaining the new scaled dot-product attention output, as depicted in Fig. 2(b). Given the implementation of residual attention, it can be easily included in both existing single-scale or multi-scale vision transformer architectures, maintaining a similar computational cost.
@@ 3.4 Attention Globality
Attention globality is a concept that reflects the capacity of the attention mechanism to encompass and consider information from all patches across an image rather than being restricted to local or nearby elements. In other words, it represents the attention’s global receptive field. This capability becomes particularly pronounced in the deeper layers of a ViT because these layers can extract high-level representations that account for long-range relationships between patches. To quantify the globality of the attention mechanism across both heads and layers, researchers in [7] introduce a metric known as the non-locality metric. This metric calculates, for each query patch $i$, the relative positional distances to all key patches $j$ weighted by their attention scores $A^{i,j}_{l,h}$. The resulting sum is then averaged by the number of patches to derive the non-locality metric for a specific head $h$. This can further be averaged across attention heads to obtain the non-locality metric for the entire layer. Mathematically, this metric is defined as follows:
$D_{l,h}=\frac{1}{N}\sum_{i,j}A^{i,j}_{l,h}||\delta^{i,j}||,\\ D_{l}=\frac{1}{H}\sum_{h}D_{l,h}.$ (11)
Here, $||\delta^{i,j}||$ represents the relative positional distance between query patch $i$ and key patch $j$, $D_{l,h}$ denotes the non-locality metric within a given layer $l$ and head $h$, while $D_{l}$ represents $D_{l,h}$ averaged along the head dimension. Moreover, since $||\delta^{i,j}||$ remains constant throughout the network (i.e., the relative positions of patches do not change), the value of $D_{l,h}$ relies entirely on $A_{l,h}$. Furthermore, as the non-locality metric $D$ increases in deeper layers (indicating a shift towards global attention), it suggests that the attention matrix $A_{l,h}$ (consequently $A_{l}$) in these layers generally has a greater capacity to capture global relationships across the different attention heads compared to attention matrix from the previous layer. This property of the attention mechanism can be harnessed to demonstrate the advantages of employing residual attention in slowing down the process of globalizing $A_{l}$. Taking into account Eq. (10), which defines the calculation of $A_{l}$ with a residual connection, it considers attention scores not only from the current layer $l$ but also from previous layers. Consequently, the globality of layer $l$ in ReViT, at a given head, can be represented as:
$D_{l,h}=\frac{1}{N}\sum_{i,j}[\alpha(A^{i,j}_{l,h})+(1-\alpha)(A^{i,j}_{l-1,h})]||\delta^{i,j}||.$ (12)
@@ 4 Experiments
This section details the implementation and evaluation of the proposed ReViT architecture. We conducted several experiments using well-established benchmark datasets for image classification, object detection, and instance segmentation to comprehensively assess the proposed architecture’s effectiveness and generalization across various computer vision tasks in line with our goals.
4.1 Datasets and Performance Evaluation
To rigorously evaluate ReViT’s performance and versatility, we conduct an extensive empirical analysis across a diverse set of image classification benchmarks. These benchmarks were chosen to encompass varying scales and complexities, ensuring a comprehensive assessment under various dataset scale scenarios. Specifically, we leverage ImageNet-1K, the de facto standard for image classification, to evaluate ReViT’s ability to learn and generalize from a vast and diverse set of over 1 million images spanning 1,000 categories. In contrast, CIFAR-10 and CIFAR-100, with their smaller size and limited classes, allow us to assess ReViT’s performance when training data is scarce, a common challenge in real-world applications. Finally, we include the fine-grained datasets, Oxford Flowers-102 and Oxford-IIIT Pets, to scrutinize ReViT’s robustness to image transformations and its capacity to discern subtle visual differences between visually similar classes. This evaluation is crucial for applications requiring precise identification.
To comprehensively assess ReViT’s performance and generalization capabilities, we evaluate it on multiple fronts. First, we benchmark ReViT against standard ViT architectures and state-of-the-art image classification models using established metrics such as top-1 and top-5 accuracy. Next, we delve into ReViT’s internal workings by quantitatively analyzing attention maps with the non-locality metric defined in Equations (11-12) and qualitatively examining feature maps using GradCAM [33]. This analysis reveals how ReViT’s attention mechanism contributes to its improved performance. Finally, we demonstrate ReViT’s versatility by integrating it into multiscale transformer architectures (Swin and MViTv2). We assess its performance on object detection and instance segmentation tasks using the COCO2017 dataset, employing mean Average Precision (mAP) on bounding boxes and produced masks as the evaluation metrics, respectively. This multifaceted evaluation provides a holistic understanding of residual attention potential across diverse visual recognition tasks.
@@ 4.2 Implementation Details
The presented ReViT model is an augmented variant of the standard ViT architecture, tailored to enhance its performance by integrating the presented residual attention module. Specifically, we integrated the residual attention module on top of the standard ViT implementation by employing the PyTorch framework. Regarding the network version used in this study, we only relied on the base version of ViT containing 12 layers and denominated as ViT-B. Consequently, the augmented ViT is referred to as ReViT-B. To verify the versatility and effectiveness of ReViT, we also integrated the proposed approach to multiscale architectures, specifically the MViTv2 and Swin transformers. These experiments consider their respective tiny versions, i.e., MViTv2-T and Swin-T, each comprising 12 layers. The choice of these smaller-scale models, referred to as ReMViTv2-T and ReSwin-T in their augmented forms, was driven by a strategic aim to reduce computational load. This is particularly crucial as larger models typically necessitate extensive pre-training on vast datasets such as ImageNet21K. By focusing on these scaled-down versions, we aim to demonstrate that the proposed enhancements can deliver performance improvements even without the extensive computational resources typically required for larger networks.
To train the augmented version of the networks, we follow the same settings as reported in each architecture’s original study using an NVIDIA V100 GPU with 32GB of RAM. Specifically, for image classification on ImageNet1K, all models were trained for $300$ epochs, with an input image size of $224\times 224$, using gradient clipping and cosine scheduling learning rate warmup/decay. For the ReViT-B network training, Adam [17] was used as optimizer with a base learning rate set to $0.001$, $30$ warmup epochs, an effective batch size of $4096$, and the weight decay set to $0.3$. Instead, for ReSwin-T model training, AdamW optimizer [28] was used by setting a base learning rate to $0.001$, $20$ warmup epochs, an effective batch size of $1024$, and the weight decay set to $0.05$. Finally, the remaining ReMViTv2-T network training was performed using the AdamW optimizer with a base learning rate set to $0.002$, $70$ warmup epochs, an effective batch size of $2048$, and the weight decay set to $0.1$. Regarding other image classification datasets, we use the same training setups reported in [6] for fair comparisons. Moreover, regarding the object detection and instance segmentation tasks, since such tasks require local awareness of features, the experiments were performed only using ReSwin-T and ReMViTv2-T. These models were trained for $36$ epochs using AdamW as an optimizer and linear scheduling learning rate decay. For ReSwin-T network training, the initial learning rate was set to $0.0001$, with a weight decay of $0.05$, and a batch size set to $16$. Instead, for ReMViTv2-T model training, the initial learning rate was set to $0.00016$, with a weight decay of $0.1$, and a batch size set to $64$. For more in-depth implementation details, we point the reader to our code repository  [^2] https://github.com/ADiko1997/ReViT.
@@ 4.3 Image classification
Subsequently, ReViT-B was tested on multiple image classification benchmarks, including CIFAR-10, CIFAR-100, Oxford Flowers-102, and Oxford-IIIT Pet, to simulate scenarios closer to real-world applications with data of small scale and low quality. Similar to ImageNet1K, top-1 accuracy on single crop images is used to assess the models. The performance of ReViT-B is then compared with that of the ViT-B model, and the results are presented in Table 3. Notably, ReViT-B, with the residual attention, consistently outperforms ViT-B across all datasets with a minimum improvement of 0.6% in CIFAR10 and a maximum of 3.4% in CIFAR100. This improvement can be attributed to the residual attention module’s ability to capture a wider range of visual features, thus enhancing the network’s capability to distinguish between different classes based on the image content. Note that the smaller improvement in CIFAR10 compared to CIFAR100 is due to the lower complexity of the former and the high accuracy of the standard ViT. The other models were not tested on these datasets because their official studies do not provide experimental results and focus only on large-scale datasets.
@@
Model Image Size CIFAR-10 CIFAR-100 Oxford-Flowers 102 Oxford-IIIT Pet
ViT-B [6] $384\times 384$ 98.1 87.1 89.5 93.8
ReViT-B $384\times 384$ 98.7 90.5 91.1 94.9
@@ 4.4 Translation Invariance
In this study, the datasets used for visual recognition typically feature the object of interest centered within the image. This is in stark contrast to real-world scenarios where objects can be situated anywhere within the scene. To assess ReViT robustness, two experiments were devised to simulate varying scales and positions of the target object within the images as shown in Fig. 5. The performance of ReViT-B and ViT-B in terms of top-1 accuracy on the Oxford Flowers-102 and Oxford-IIIT Pet datasets are compared in these experiments.
The first experiment evaluates the models’ ability to handle horizontal and vertical translations. To achieve this, the input images are horizontally and vertically shifted on the image plane by {15, 30, 45, 60}of pixels, with the empty areas filled with zeros to maintain the same input image size as shown in Fig. 5(a) and (b). As indicated in Table 4 and Fig. 6(a), both models exhibit a similar and impressive level of horizontal translation invariance, with only a minor reduction in performance. Overall, ReViT-B demonstrates slightly superior horizontal translation invariance, especially at the 60% translation level. For the Oxford Flowers-102 dataset, ReViT-B performance decreases by 2.9% less than ViT-B at a 60% translation, while for the Oxford-IIIT Pet dataset, the decrease is 1.0% less. Similar trends are observed for vertical translation invariance, as illustrated in Table 5 and Fig. 6(b). Although both models show a comparable drop in performance, ReViT-B displays better overall vertical translation invariance, particularly on the Oxford Flowers-102 dataset, where its performance decreases by 3.6% less than ViT-B.
In the second experiment, the scale invariance of both models is evaluated, and the results are presented in Table 6 and Fig. 6(c). To generate images at different scales, the original images are resized by reducing their width and height by {15, 30, 45, 60}%, with zero-padding to maintain the correct input size as illustrated in Fig. 5(c). Unlike the first experiment, ReViT-B consistently outperforms ViT-B in terms of scale invariance. Most notably, on the Oxford Flowers-102. Particularly, this dataset exhibits high intra-class variation and inter-class similarity for certain categories. Thus, accurately identifying flower categories requires capturing fine details like shape, color, and proportions. When images are downscaled significantly (beyond 45%), the details become much smaller and the number of informative tokens in the sequence decreases. Additionally, the effective distance between informative tokens shrinks as patched regions (filled with zeros) do not affect the attention calculations (i.e., multiplication with 0, no information to aggregate). This reduced distance allows attention to rapidly reach global peaks, leading to earlier feature collapse. Under these conditions, the vanilla ViT, lacking an anti-collapse mechanism, experiences a significant performance drop of up to 58.5 percentage points. This is 19.5 percentage points or 33.3% more than ReViT’s reduction. The key difference lies in ReViT’s residual attention mechanism, which counteracts feature collapsing. These results highlight that ReViT-B possesses significantly better scale invariance compared to ViT-B.
@@ 4.5 Residual Attention and Globality
The investigation into the global reach of MHSA within vision transformers, both with and without residual attention, aims to examine the hypothesis that residual attention decelerates the globalization of the attention mechanism. To assess this hypothesis quantitatively, we employ the non-locality metric calculated according to Eq. (11). Additionally, the metric is computed at both the head and layer levels. The results of this analysis, depicted in Fig. 7, display the average metric across 256 input images from the ImageNet1K validation set for ViT-B and ReViT-B models trained on the same dataset. The x-axis on the plots corresponds to the layer index, while the mean non-locality for each head or layer is represented on the y-axis. Fig. 7(a) illustrates that in ViT-B, the non-locality of each head increases as it delves deeper into the layers, approaching globality which corresponds to the findings in [6, 7]. A similar trend is observed in ReViT-B, as seen in Fig. 7(b), albeit with some noteworthy distinctions. Notably, ViT-B exhibits a more consistent and monotonic increase in non-locality, with minimal variations between different heads, particularly after layer 5. In contrast, ReViT-B displays more diverse differences between heads, with some heads finding difficulty in escaping the locality while others become completely global, resulting in a less monotonic increase in non-locality. Such behavior is caused by the residual attention mechanism, which induces locality from previous layers into the attention of current layers. In summary, the results substantiate the hypothesis that the attention receptive field of heads in vision transformers becomes global in deeper layers, a phenomenon observed in both ViT-B and ReViT-B models. However, these findings also reveal that ReViT-B retains the ability to extract global relationships in deeper layers while still maintaining some heads that remain localized with relatively low non-locality values. Simultaneously, ViT-B is confined solely to global relationships, leading to feature collapse. To provide a broader perspective, Fig. 7(c) presents the non-locality of ViT-B and ReViT-B at the layer level, represented as the average non-locality metric across all heads within each layer. It is evident that when computed using the same 256 input examples, ReViT-B consistently maintains lower non-locality values (except the first), indicating its ability to incorporate low-level features, with the help of relatively local heads, into the representations learned from heads that exhibit non-locality escape (i.e., become global).